OpenStudy (anonymous):

Solve. 8 = 2^(x + 4)

OpenStudy (precal):

log both sides

OpenStudy (precal):

drop the power (x+4) in front of the log

OpenStudy (anonymous):

x = –4 x = –1 x = 0 x = 7

OpenStudy (precal):

I don't need the choices, I am talking you through the process. Process is more important since it always leads to the correct solution

OpenStudy (precal):

You will need this skill if you plan to attend college, this a necessary skill

OpenStudy (anonymous):

2^3 = 2^(x+4) So 3 = x +4